Calculate and sign the tax multiplier (dY/dt) in the following model. Y = C[(1-t)Y] + I[ i ] + G M = L[i, (1-t)Y]
Answer #1

Here, Y = C[(1-t)Y] + I[i] + G

Taking a total differential of Y

dY = C/T[dY - tdY] + dG ...... Since I is not a function of T (income is not a function of taxes), it's derivative will be zero

dY = bdY - btdY + dG ..... Where b = C/T

[ 1- b + bt] dY = (1-b) dG

dY/dG = 1-b/ (1-b+bt)

dY/dT = 1-b/ (1-b+bt) where b = C/T

The tax multiplier always has a negative sign becuase as tax decreases, income increases